### v1

# Fourier growth of degree $2$ polynomials

Abstract.
We prove bounds for the absolute sum of all level-$k$ Fourier coefficients for $(-1)^{p(x)}$, where polynomial $p:\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}$ is of degree $\leq 2$. First we prove the bound $\lesssim(1+\sqrt{2})^{k}$. Then, by a different method, we make it sharp: $\lesssim k^{-1/2}(1+\sqrt{2})^{k}$.
Key words and phrases:
Pseudorandom generators, Fourier weight of $\mathbf{F}_{2}$ polynomials, vectors of a fixed boolean weight
2010 Mathematics Subject Classification:
68Q17; 46B10; 60E15
LB, JS, HZ, AV:This work was completed while the authors were in residence at the Hausdorff Research Institute for Mathematics in Bonn, during the trimester program “Boolean Analysis in Computer Science”, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813; L.B. is supported by the Collaborative Research Center 1060 funded by the Deutsche Forschungsgemeinschaft and the Hausdorff Center for Mathematics, funded by the DFG under Germany’s Excellence Strategy - GZ 2047/1. O.K. was supported in part by a grant from the Israel Science Foundation (ISF Grant No. 1774/20), and by a grant from the US-Israel Binational Science Foundation and the US National Science Foundation (BSF-NSF Grant No. 2020643. J.S. is supported by Chris Umans’ Simons Institute Investigator Grant. A.V. is supported by NSF DMS-2154402 and by Hausdorff Center for Mathematics.
1. Introduction
We want to estimate independently of $n$ the $k$-th Fourier weight of $(-1)^{p(x)}$, where $p:\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}$ a polynomial of degree $\leq d$. Here, the $k$-th Fourier weight of a function $f:\mathbf{F}_{2}^{n}\to\\ mathbf{C}$ refers to the quantity
$\sum_{|A|=k}|\hat{f}(A)|\,,$
where
$\hat{f}(A)=\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}f(x)(-1)^{x_{A}}\,,\qquad x_{A}=\sum_{i\in A}x_{i}\,.$
For polynomials $p$ of degree 00 or $1$, the absolute sum of all Fourier coefficients of $(-1)^{p(x)}$ is exactly $1$. In particular the $k$-th Fourier weight is at most $1$ for all $k$.
For polynomials of degree at most $2$, we will prove the following.
Theorem 1.
For any degree $2$ polynomial $p:\mathbf{F}^{n}_{2}\to\mathbf{F}_{2}$, the $k$-th Fourier weight of $(-1)^{p(x)}$ is at most
$(1+\sqrt{2})^{k}\,.$
This confirms a conjecture of Chattopadhyay, Hatami, Hosseini and Lovett, see [CHHL, Open Problem 1.7], for polynomials of degree $2$. They state in [CHHL] that they have a proof for degree $2$ polynomomials. Pooya Hatami has kindly sketched to us what argument they had in mind, we reproduce it in Section 3. Up to a factor $k^{-1/2}$, it gives the same bound.
2. Proof of Theorem 1
We denote by $W(2,k)$ the maximum over all polynomials of degree $2$ on $\mathbf{F}_{2}^{n}$ of the $k$-th Fourier weight of $(-1)^{p(x)}$:
$W(2,k)=\max_{\begin{subarray}{c}n\geq 1\\ \mathrm{deg}\,p\leq 2\end{subarray}}\sum_{A\subset[n],|A|=k}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|\,.$ (2.1)
Here we abbreviate $x_{A}=\sum_{i\in A}x_{i}$. Theorem 1 claims that
$W(2,k)\leq(1+\sqrt{2})^{k}\,.$ (2.2)
We will prove this by induction on the number of variables $n$ and on $k$. Note that the estimate is obvious when $k\leq 0$ or $n<k$, since then we have $W(2,k)=0$. It also holds for $n=k$, since then $W(2,k)\leq 1$. So let $k\geq 1$ and $n\geq k+1$.
We denote the maximum $k$-th Fourier weight among all $n$ variable degree $2$ polynomials by $W(2,k,n)$:
$W(2,k,n)=\max_{\mathrm{deg}\,p\leq 2}\sum_{A\subset[n],|A|=k}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|\,.$
For our induction step, we assume that for all $k^{\prime}<k$
$W(2,k^{\prime})\leq(1+\sqrt{2})^{k^{\prime}}$ (2.3)
and for all $n^{\prime}<n$
$W(2,k,n^{\prime})\leq(1+\sqrt{2})^{k}\,.$ (2.4)
Our goal is to deduce from this that
$W(2,k,n)\leq(1+\sqrt{2})^{k}\,,$
this will complete the proof of the theorem.
Let $p$ be a maximizing polynomial of degree at most $2$ in $n$ variables, such that
$W(2,k,n)$ $=\sum_{A\subset[n],|A|=k}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|$
$=\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}\sum_{A\subset[n],|A|=k}\varepsilon_{A}\cdot(-1)^{x_{A}}$ (2.5)
for certain $\varepsilon_{A}\in\{-1,1\}$.
If $p$ is affine then we have
$p(x)=x_{1}+\dotsc+x_{m}+const$
for some $m$. In that case (2.5) is $1$ if $k=m$ and 00 if $k\neq m$, so $W(2,k,n)\leq 1$.
If $p$ is not affine, then we can write it (perhaps after permuting the variables) as
$p(x)=p_{0}(\bar{x})+x_{1}p_{1}(\bar{x})+x_{2}p_{2}(\bar{x})+x_{1}x_{2}\,,$
where $\bar{x}=(x_{3},\dotsc,x_{n})$. For a set $A\subset[n]$, we abbreviate $\bar{A}=A\cap\{3,\dotsc,n\}=A\setminus\{1,2\}$. Writing out the sum over $x_{1},x_{2}$ in (2.5), we then obtain
$W(2,k,n)$
$=\frac{1}{4}\frac{1}{2^{n-2}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})}\cdot$
$\Big{[}\sum_{A\cap\{1,2\}=\emptyset}\varepsilon_{A}(-1)^{x_{\bar{A}}}+\sum_{A\cap\{1,2\}=\{1\}}\varepsilon_{A}(-1)^{x_{\bar{A}}}+\sum_{A\cap\{1,2\}=\{2\}}\varepsilon_{A}(-1)^{x_{\bar{A}}}+\sum_{\{1,2\}\subset A}\varepsilon_{A}(-1)^{x_{\bar{A}}}\Big{]}$
$+\frac{1}{4}\frac{1}{2^{n-2}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})}(-1)^{p_{1}(\bar{x})}\cdot$
$\Big{[}\sum_{A\cap\{1,2\}=\emptyset}\varepsilon_{A}(-1)^{x_{\bar{A}}}-\sum_{A\cap\{1,2\}=\{1\}}\varepsilon_{A}(-1)^{x_{\bar{A}}}+\sum_{A\cap\{1,2\}=\{2\}}\varepsilon_{A}(-1)^{x_{\bar{A}}}-\sum_{\{1,2\}\subset A}\varepsilon_{A}(-1)^{x_{\bar{A}}}\Big{]}$
$+\frac{1}{4}\frac{1}{2^{n-2}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})}(-1)^{p_{2}(\bar{x})}\cdot$
$\Big{[}\sum_{A\cap\{1,2\}=\emptyset}\varepsilon_{A}(-1)^{x_{\bar{A}}}+\sum_{A\cap\{1,2\}=\{1\}}\varepsilon_{A}(-1)^{x_{\bar{A}}}-\sum_{A\cap\{1,2\}=\{2\}}\varepsilon_{A}(-1)^{x_{\bar{A}}}-\sum_{\{1,2\}\subset A}\varepsilon_{A}(-1)^{x_{\bar{A}}}\Big{]}$
$-\frac{1}{4}\frac{1}{2^{n-2}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})}(-1)^{p_{1}(\bar{x})}(-1)^{p_{2}(\bar{x})}\cdot$
$\Big{[}\sum_{A\cap\{1,2\}=\emptyset}\varepsilon_{A}(-1)^{x_{\bar{A}}}-\sum_{A\cap\{1,2\}=\{1\}}\varepsilon_{A}(-1)^{x_{\bar{A}}}-\sum_{A\cap\{1,2\}=\{2\}}\varepsilon_{A}(-1)^{x_{\bar{A}}}+\sum_{\{1,2\}\subset A}\varepsilon_{A}(-1)^{x_{\bar{A}}}\Big{]}.$
We will simplify this using the following identity, which is valid for all $p_{1},p_{2}\in\mathbf{F}_{2}$:
$1+(-1)^{p_{1}}+(-1)^{p_{2}}-(-1)^{p_{1}+p_{2}}=2(-1)^{p_{1}p_{2}}\,.$
Replacing $p_{1}$ by $p_{1}+1$ and $p_{2}$ by $p_{2}+1$ yields also
$1-(-1)^{p_{1}}+(-1)^{p_{2}}+(-1)^{p_{1}+p_{2}}=2(-1)^{p_{1}p_{2}+p_{2}}\,,$
$1+(-1)^{p_{1}}-(-1)^{p_{2}}+(-1)^{p_{1}+p_{2}}=2(-1)^{p_{1}p_{2}+p_{1}}\,,$
and
$1-(-1)^{p_{1}}-(-1)^{p_{2}}-(-1)^{p_{1}+p_{2}}=2(-1)^{p_{1}p_{2}+p_{1}+p_{2}}\,.$
Using the first identity on the first terms in square brackets, the second identity on the second terms etc. in the expansion of $W(2,k,n)$ yields that $W(2,k,n)$ equals
$\frac{1}{2}\frac{1}{2^{n-1}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x})}\sum_{A\subset[n]\setminus\{1,2\},|A|=k}\varepsilon_{A}(-1)^{x_{A}}$
$+\frac{1}{2}\frac{1}{2^{n-1}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x})+p_{2}(\bar{x})}\sum_{A\subset[n]\setminus\{1,2\},|A|=k-1}\varepsilon_{A\cup\{1\}}(-1)^{x_{A}}$
$+\frac{1}{2}\frac{1}{2^{n-1}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x})+p_{1}(\bar{x})}\sum_{A\subset[n]\setminus\{1,2\},|A|=k-1}\varepsilon_{A\cup{2}}(-1)^{x_{A}}$
$+\frac{1}{2}\frac{1}{2^{n-1}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}-(-1)^{p_{0}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x})+p_{1}(\bar{x})+p_{2}(\bar{x})}\sum_{A\subset[n]\setminus\{1,2\},|A|=k-2}\varepsilon_{A\cup\{1,2\}}(-1)^{x_{A}}\,.$
The expression in the first line, for fixed $A$, is the $A$-th Fourier coefficient of the function $(-1)^{q_{1}(\bar{x})}$, where
$q_{1}(\bar{x})=p_{0}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x}).$
This $q_{1}$ is a degree $2$ polynomial, because $p_{1}$ and $p_{2}$ are by assumption affine polynomials. Furthermore, $q_{1}$ only depends on $n-2$ variables. Hence, the absolute value of first line is at most $\frac{1}{2}W(2,k,n-2)$.
Similarly, the second and third line are in absolute value at most $\frac{1}{2}W(2,k-1,n-2)$, and the last line at most $\frac{1}{2}W(2,k-2,n-2)$. It follows that
$W(2,k,n)$ $\leq\frac{1}{2}W(2,k,n-2)+W(2,k-1)+\frac{1}{2}W(2,k-2)$
$\leq(\frac{1}{2}(1+\sqrt{2})^{k}+(1+\sqrt{2})^{k-1}+\frac{1}{2}(1+\sqrt{2})^{k-2})$
$=(1+\sqrt{2})^{k}\,.$
Here we used the induction hypotheses (2.3) and (2.4). This completes the proof.
3. Another approach via the structure theorem of Dickson
Pooya Hatami kindly sketched to us what proof of the estimate of Fourier weight-level-$k$ for quadratic polynomials they had in mind in [CHHL]. As [CHHL] does not have the proof of this kind of estimate we include here that proof as it seems interesting to compare it with the proof in Section 2.
It is based on a beautiful structure theorem of Dickson (see Theorem 4 Chapter 15 of [McWS]), which we formulate now in a convenient form. First notice that any quadratic polynomial $q:\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}$ can be written in the form
$q(x)=x^{t}Ax+(\ell,x)+c,$
where $A$ is an upper triangular matrix with $\mathbf{F}_{2}$ entries, $\ell$ is a binary vector and $c\in\mathbf{F}_{2}$.
Theorem 2 ([McWS] and Lemma 2.4 [BGGM]).
Let $q$ be as above and let $A$ be an upper triangular binary matrix of size $n\times n$. Let $m$ be the rank of $A+A^{t}$. Then there exists a non-singular linear mapping $T:\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}^{n}$ such that
$q\circ T^{-1}(y)=\sum_{j=1}^{m}y_{2j-1}y_{2j}+(L,y)+b,$ (3.1)
where $L\in\mathbf{F}_{2}^{n},b\in\mathbf{F}_{2}$. Moreover, if $(L,y)$ is linearly dependent on $\{y_{1},\dots,y_{2m}\}$, there exists another affine change of variable $\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}^{n}$ that transfers $q\circ T(y)$ into $Q(z),z=(z_{1},\dots,z_{n})$, where
$Q(z)=\sum_{j=1}^{m}z_{2j-1}z_{2j}+d,$ (3.2)
where $d\in\mathbf{F}_{2}$.
Those change of variables do not change the expectation over $\mathbf{F}_{2}^{n}$. One simple consequence then is the fact about the Fourier coefficients of $f:=(-1)^{q(x)}$: Namely, let $S\subset[n]$, then
$|\hat{f}(S)|$ $=|\mathbf{E}(-1)^{q(x)+\sum_{k\in S}x_{k}}|$
=$=\begin{cases}2^{-m},{\sum_{k\in S}x_{k}}\in-(L,T(x))+\text{span}\{(T(x))_{j}\}_{j=1}^{2m},\\ 0,\quad\text{otherwise}\,.\end{cases}$ (3.3)
Let us estimate the number of $S$, $|S|=k$, such that the first line holds.
3.1. Vectors of boolean weight $k$ in affine subspace of $\mathbf{F}_{2}^{n}$.
We need to prove that any affine subspace of $\mathbf{F}_{2}^{n}$ of dimension $2m$ contains at most (${2m+1\choose k}$ vectors of boolean weight $k$.
Lemma 3.
Let $W\subset\mathbb{F}_{2}^{n}$ be an affine subspace of dimension $h$. Then $W$ contains at most (${h+1\choose k}$ vectors of weight $k$.
Proof.
We write $W=w+V$ for a subspace $V$ of dimension $h$. By permuting the coordinates and applying Gaussian elimination, we can choose a basis $v^{1},\dotsc,v^{h}$ that looks as follows:
$\begin{pmatrix}--v^{1}--\\ --v^{2}--\\ \vdots\\ --v^{h}--\end{pmatrix}=\begin{pmatrix}1&0&\ldots&0&*&*&\dotsb&*\\ 0&1&\ldots&0&*&*&\dotsb&*\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&&\vdots\\ 0&0&\ldots&1&*&*&\dotsb&*\end{pmatrix}\,.$
Now we can obviously further assume that $w_{1}=\dotsb=w_{h}=0$. We will now prove the theorem by induction.
The theorem holds when $k=1$. Indeed, any sum of at least $2$ vectors $v^{j}$ is nonzero in at least two of the coordinates $1,\dotsc,h$, and hence its sum with $w$ does not have weight $1$. So there are at most $h+1$ vectors in $w+V$ that could have weight $1$, namely
$w,w+v^{1},\dotsc,w+v^{h}\,.$
Now let us assume that it holds for $k-1$. If the coefficient of $v^{1}$ in a weight $k$-vector $v\in w+V$ is nonzero, then we can write
$v=\delta^{1}+(w+v^{1}-\delta^{1})+r$
[…]
(${h\choose k-1}$
choices for $v$ with nonzero coefficient of $v^{1}$.
On the other hand, if the coefficient of $v^{1}$ vanishes, then $v$ itself is a weight $k$ vector in the dimension $h-1$ affine subspace $w+V^{\prime}=w+\langle v^{2},\dotsc,v^{h}\rangle$. By induction again, there are at most
(${h\choose k}$
such vectors $v$. Thus, in total there are at most
(${h\choose k-1}+{h\choose k}={h+1\choose k}$
vectors $v$. This completes the proof. ∎
The estimate in Lemma 3 is sharp: If $W$ is the affine subspace in $\mathbf{F}_{2}^{n}$ generated by
$\{x\in\mathbb{F}_{2}^{n}\,:\,\sum_{i=1}^{n}x_{i}=k\}$
then dimension $W$ is $n-1$, and $W$ contains exactly
(${n\choose k}={\dim W+1\choose k}$
vectors of weight $k$, namely all such vectors in $\mathbb{F}_{2}^{n}$.
3.2. Estimating the Fourier weight
Combining (3.3) and Lemma 3 shows that
$\sum_{|S|=k}|\hat{f}(S)|\leq 2^{-m}{2m+1\choose k}\,.$
To further estimate this we start with
Lemma 4.
For all $\alpha\in[0,1]$, we have
$\alpha^{\alpha}(2-\alpha)^{2-\alpha}\geq 2(\sqrt{2}-1)^{\alpha}\,.$
Proof.
Substituting $\alpha=2(1-y)$ turns this into the equivalent inequality
$\left(\frac{y}{2^{-1/2}}\right)^{y}\left(\frac{1-y}{1-2^{-1/2}}\right)^{1-y}\geq 1\,.$
It is readily verified that equality holds at $y=2^{-1/2}$. Furthermore, if $h(y)=\exp(g(y))$ denotes the left hand side of the inequality, then
$\frac{\mathrm{d}}{\mathrm{d}y}h(y)=\exp(g(y))g^{\prime}(y)=\exp(g(y))\left(\log\frac{y}{2^{-1/2}}-\log\frac{1-y}{1-2^{-1/2}}\right)\,.$
Thus, the derivative vanishes at $y=2^{-1/2}$. Finally, we have
$\frac{\mathrm{d}^{2}}{\mathrm{d}y^{2}}h(y)=\exp(g(y))((g^{\prime}(y))^{2}+g^{\prime\prime}(y))\geq\exp(g(y))g^{\prime\prime}(y)$
$=\exp(g(y))\left(\frac{1}{y}+\frac{1}{1-y}\right)\,,$
so $h$ is convex and the inequality follows. ∎
Lemma 5.
There exists a constant $C>0$ such that for all $m\geq 1$ and $0\leq k\leq m$
(${2m+1\choose k}\leq C2^{m}k^{-1/2}(1+\sqrt{2})^{k}\,.$
The number $1+\sqrt{2}$ on the right hand side cannot be replaced by a smaller number.
Proof.
From Stirling’s formula, we have for $k=\alpha(m+1/2)$:
$2^{-m}{2m+1\choose k}$
$\sim 2^{-m}\frac{1}{\sqrt{2\pi k(2-\alpha)}}\frac{(2m+1)^{2m+1}}{(\alpha(m+1/2))^{\alpha(m+1/2)}((2-\alpha)(m+1/2))^{(2-\alpha)(m+1/2)}}$
$=\frac{1}{\sqrt{\pi k(2-\alpha)}}\left(\frac{2}{\alpha^{\alpha}(2-\alpha)^{(2-\alpha)}}\right)^{m+1/2}\,.$
Setting $\alpha=2-\sqrt{2}$ (which is the point where the inequality of the previous lemma is sharp), this equals
$Ck^{-1/2}(1+\sqrt{2})^{\alpha(m+1/2)}=Ck^{-1/2}(1+\sqrt{2})^{k}\,.$
This shows sharpness of the number $1+\sqrt{2}$.
To obtain the claimed inequality and not just an asymptotic bound for large $m$ and $k$, we use the following bound in place of Stirling’s formula:
$\sqrt{2\pi n}\left(\frac{n}{e}\right)^{n}e^{\frac{1}{12n}-\frac{1}{360n^{3}}}\leq n!\leq\sqrt{2\pi n}\left(\frac{n}{e}\right)^{n}e^{\frac{1}{12n}}\,.$
This implies
$2^{-m}{2m+1\choose k}$
$<\frac{1}{\sqrt{2\pi\alpha(2-\alpha)}}\left(\frac{2}{\alpha^{\alpha}(2-\alpha)^{(2-\alpha)}}\right)^{m+1/2}e^{\frac{1}{24m}-\frac{1}{12k}-\frac{1}{12(2m-k)}+\frac{1}{360(2m-k)^{3}}+\frac{1}{360k^{3}}}$
$\leq\frac{e^{3}}{\sqrt{2\pi k}}\left(\frac{2}{\alpha^{\alpha}(2-\alpha)^{(2-\alpha)}}\right)^{m+1/2}\,.$
Using Lemma 4
$\leq Ck^{-1/2}(1+\sqrt{2})^{k}\,.$
We got the desired estimate for $k\leq m$. If $m+1\leq k\leq 2m+1$ we just put $\ell=2m+1-k<k$ and get the previous estimate with $\ell$ instead of $k$. We are done:
$W(2,k)\leq Ck^{-1/2}(1+\sqrt{2})^{k},$
where we used that function $x\to\frac{(1+\sqrt{2})^{x}}{x^{1/2}}$ increases on $[1,\infty)$.
3.3. Sharpness
In fact, the arguments in this section were all sharp: Let
$q(x)=\sum_{j=1}^{m}x_{2j}x_{2j-1}$
and let
$k=(2-\sqrt{2})m\,.$
We use (3.3), and note that we are in the equality case of both Lemma 3 and Lemma 5, to obtain in this case
$\sum_{|S|=k}|\hat{f}(S)|\sim Ck^{-1/2}(1+\sqrt{2})^{k}\,.$
Thus the bound in Theorem 1 is sharp, up to the factor $k^{-1/2}$.