### v2

# Fourier growth of degree $2$ polynomials

Abstract.
We prove bounds for the absolute sum of all level-$k$ Fourier coefficients for $(-1)^{p(x)}$, where polynomial $p:\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}$ is of degree $1$ or degree $2$.
Key words and phrases:
Pseudorandom generators, Fourier weight of $\mathbf{F}_{2}$ polynomials, vectors of a fixed boolean weight
2010 Mathematics Subject Classification:
68Q17; 46B10; 60E15
This work was completed while the authors were in residence at the Hausdorff Research Institute for Mathematics in Bonn, during the trimester program “Boolean Analysis in Computer Science”, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813;
L.B. is supported by the Collaborative Research Center 1060 funded by the Deutsche Forschungsgemeinschaft;
J.S. is supported by Chris Umans’ Simons Institute Investigator Grant;
A.V. is supported by NSF DMS-2154402 and by the Hausdorff Center for Mathematics;
H.Z. is supported by NSF DMS-2453408.
1. Introduction
Let $\mathbf{F}_{2}=\{0,1\}.$ Any function $f:\mathbf{F}_{2}^{n}\to\mathbf{C}$ has the Fourier–Walsh expansion
$f(x)=\sum_{A\subset[n]}\widehat{f}(A)(-1)^{x_{A}},$ (1.1)
where $A\subset[n]:=\{1,2,\dots,n\}$ and $x_{A}=\sum_{i\in A}x_{i}$. The degree of $f$ is defined as $\deg(f):=\max_{\widehat{f}(A)\neq 0}|A|$, where $|A|$ denotes the cardinality of $A$.
Boolean functions $f:\mathbf{F}_{2}^{n}\to\{-1,1\}$ play an important role in theoretical computer science and social choice theory. Of particular interest to us are those of the form $f(x)=(-1)^{p(x)}$ with $p:\mathbf{F}^{n}_{2}\to\mathbf{F}_{2}$ a polynomial of small degree.
We want to estimate independently of $n$ the $k$-th Fourier weight of $(-1)^{p(x)}$, where $p:\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}$ a polynomial of degree at most $d$. Here, the $k$-th Fourier weight of a function $f:\mathbf{F}_{2}^{n}\to\mathbf{C}$ refers to the quantity
$\sum_{|A|=k}|\widehat{f}(A)|\,$,
where
$\widehat{f}(A)=\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}f(x)(-1)^{x_{A}}\,.$
In fact, Chattopadhyay, Hatami, Hosseini and Lovett [CHHL18, Open problem 1.7] conjectured the following.
Conjecture 1.
Let $f(x)=(-1)^{p(x)},x\in\mathbf{F}_{2}^{n}$ with $p:\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}$ being a polynomial of degree at most $d$. Then there exists a constant $c_{d}>0$ depending only $d$ such that
$\sum_{A:|A|=k}|\widehat{f}(A)|\leq c_{d}^{k}$
for all $k\geq 1$. In particular, the constant $c_{d}=2^{O(d)}$ should work.
As explained in [CHHL18], it is necessary to assume the exponential dependence on $k$. In case $\deg(p)\leq 1$, we may write $p(x)=a+x_{i_{1}}+\cdots+x_{i_{m}}$ for $1\leq i_{1}<\dots<i_{m}\leq n$ and some $a\in\mathbf{F}_{2}$. Then
$|\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)+x_{A}}|=\begin{cases}2^{n}&A=\{i_{1},\dots,i_{m}\}\\ 0.&\textnormal{otherwise}\end{cases}$
Therefore, we have
$\sum_{A\subset[n]}|\widehat{f}(A)|=\frac{1}{2^{n}}\sum_{A\subset[n]}|\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)+x_{A}}|=1$
and thus
$\sum_{A:|A|=k}|\widehat{f}(A)|\leq 1.$
For $p$ of degree at most 2, we will prove by a simple induction argument the following.
Theorem 1.
For any polynomial $p:\mathbf{F}^{n}_{2}\to\mathbf{F}_{2}$ of degree at most $2$, the $k$-th Fourier weight of $f(x)=(-1)^{p(x)}$ is at most $(1+\sqrt{2})^{k}$, that is,
$\sum_{A:|A|=k}|\widehat{f}(A)|\leq(1+\sqrt{2})^{k}.$
This confirms a conjecture in [CHHL18] for polynomials of degree $2$. We note that the authors state in [CHHL18] to already have a proof in that case, which appears however not to be published. In Section 3, we give another proof of Theorem 1 based on the strategy they had in mind. We thank Pooya Hatami for kindly sharing it with us. It gives the following sharp upper bound.
Theorem 2.
For any polynomial $p:\mathbf{F}^{n}_{2}\to\mathbf{F}_{2}$ of degree at most $2$, the $k$-th Fourier weight of $f(x)=(-1)^{p(x)}$ is
$\sum_{A:|A|=k}|\widehat{f}(A)|\leq\sqrt{\frac{2}{\pi}}k^{-1/2}(1+\sqrt{2})^{k}.$
This is sharp in the sense that there exist pairs $(p,k)$ for which the ratio of right hand side and left hand side gets arbitrarily close to $1$.
Pseudorandom generators and fractional pseudorandom generators are constructed in [CHHL18], Section 4, given that the exponential bounds on Fourier weight of $(-1)^{p}$ for high degree polynomials $p$ are proved.
Pseudorandom generators are widely studied in complexity theory. There are several general frameworks used to approach their possible construction. The literature is vast, one can read a survey [Gol10] for example. A harmonic analysis approach was suggested in [CHHL18], and it requires the good estimate of Fourier weights. The reader can see other approaches in e. g. [NW94], [HLW06].
In [CGLLS20] it was shown that for polynomial error, Fourier bounds on the first $O(\log n)$ levels is sufficient to recover the seed length (the smallness of pseudorandom generator) obtained in [CHLT19], which required bounds on the entire tail.
2. Proof of Theorem 1 by induction
We denote by $W(2,k,n)$ the maximum over all polynomials of degree $2$ on $\mathbf{F}_{2}^{n}$ of the $k$-th Fourier weight of $(-1)^{p(x)}$:
$W(2,k,n)=\max_{\mathrm{deg}\,p\leq 2}\sum_{A\subset[n],|A|=k}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|\,.$ (2.1)
Here we abbreviate $x_{A}=\sum_{i\in A}x_{i}$. Theorem 1 claims that for all $0\leq k\leq n$
$W(2,k,n)\leq(1+\sqrt{2})^{k}\,.$ (2.2)
We will prove this by induction on the number of variables $n$ and on $k$. Note that the estimate is obvious when $k=0$ or $k=n$, since then we have $W(2,k,n)\leq 1$. So let $k\geq 1$ and $n\geq k+1$.
For our induction step, we assume that for all $k^{\prime}<k$ and all $n^{\prime}$
$W(2,k^{\prime},n^{\prime})\leq(1+\sqrt{2})^{k^{\prime}}$ (2.3)
and for all $n^{\prime}<n$
$W(2,k,n^{\prime})\leq(1+\sqrt{2})^{k}\,.$ (2.4)
Our goal is to deduce from this that
$W(2,k,n)\leq(1+\sqrt{2})^{k}\,$,
this will complete the proof of the theorem.
Let $p$ be a maximizing polynomial of degree at most $2$ in $n$ variables, such that
$W(2,k,n)$ $=\sum_{A\subset[n],|A|=k}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|.$ (2.5)
If $p$ is affine then we have
$p(x)=a+x_{i_{1}}+\dotsb+x_{i_{m}}$
for some $m$ and $a\in\mathbf{F}_{2}$. In that case, as explained above, (2.5) is $1$ if $k=m$ and 0 if $k\neq m$, so $W(2,k,n)\leq 1$.
If $p$ is not affine, then we write it, perhaps after permuting the variables, as
$p(x)=p_{0}(\bar{x})+x_{1}p_{1}(\bar{x})+x_{2}p_{2}(\bar{x})+x_{1}x_{2}\,$, (2.6)
where $\bar{x}=(x_{3},\dotsc,x_{n})$. For a set $A\subset[n]$, we abbreviate $\bar{A}=A\setminus[2]=A\cap\{3,\dotsc,n\}$. According to our notation, $\bar{x}_{\bar{A}}=\sum_{i\in[n]\setminus[2]}x_{i}$. Writing out the sum over $x_{1},x_{2}$ in (2.5), we then obtain
$W(2,k,n)=I_{0}+I_{1}+I_{2}+I_{12}$
where (we omit the constraints $A\subset[n]$ and $|A|=k$ for notational convenience)
$I_{0}$ $=\sum_{A\cap\{1,2\}=\emptyset}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|,$
$I_{1}$ $=\sum_{A\cap\{1,2\}=\{1\}}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|,$
$I_{2}$ $=\sum_{A\cap\{1,2\}=\{2\}}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|,$
$I_{12}$ $=\sum_{A\cap\{1,2\}=\{1,2\}}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|.$
Recall that by (2.6)
$p(0,0,\bar{x})=p_{0}(\bar{x}),\qquad p(0,1,\bar{x})=p_{0}(\bar{x})+p_{2}(\bar{x})$
and
$p(1,0,\bar{x})=p_{0}(\bar{x})+p_{1}(\bar{x}),\qquad p(1,1,\bar{x})=1+p_{0}(\bar{x})+p_{1}(\bar{x})+p_{2}(\bar{x}).$
Consider the term $I_{0}$ first. Since $A\cap\{1,2\}=\emptyset$, we have $(-1)^{x_{A}}=(-1)^{\bar{x}_{\bar{A}}},$ and thus
$I_{0}=\frac{1}{2^{n}}\sum_{A\cap\{1,2\}=\emptyset}\left|\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{\bar{x}_{\bar{A}}+p_{0}(\bar{x})}\left(1+(-1)^{p_{1}(\bar{x})}+(-1)^{p_{2}(\bar{x})}-(-1)^{p_{1}(\bar{x})+p_{2}(\bar{x})}\right)\right|.$
The other three terms can be rewritten in a similar way. For $I_{1}$, since $A\cap\{1,2\}=\{1\}$, we have $(-1)^{x_{A}}=(-1)^{x_{1}+\bar{x}_{\bar{A}}},$ and thus
$I_{1}=\frac{1}{2^{n}}\sum_{A\cap\{1,2\}=\{1\}}\left|\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{\bar{x}_{\bar{A}}+p_{0}(\bar{x})}\left(1-(-1)^{p_{1}(\bar{x})}+(-1)^{p_{2}(\bar{x})}+(-1)^{p_{1}(\bar{x})+p_{2}(\bar{x})}\right)\right|.$
For $I_{2}$, since $A\cap\{1,2\}=\{2\}$, we have $(-1)^{x_{A}}=(-1)^{x_{2}+\bar{x}_{\bar{A}}},$ and thus
$I_{2}=\frac{1}{2^{n}}\sum_{A\cap\{1,2\}=\{2\}}\left|\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{\bar{x}_{\bar{A}}+p_{0}(\bar{x})}\left(1+(-1)^{p_{1}(\bar{x})}-(-1)^{p_{2}(\bar{x})}+(-1)^{p_{1}(\bar{x})+p_{2}(\bar{x})}\right)\right|.$
For $I_{12}$, since $A\cap\{1,2\}=\{1,2\}$, we have $(-1)^{x_{A}}=(-1)^{x_{1}+x_{2}+\bar{x}_{\bar{A}}},$ and thus
$I_{12}=\frac{1}{2^{n}}\sum_{A\cap\{1,2\}=\{1,2\}}\left|\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{\bar{x}_{\bar{A}}+p_{0}(\bar{x})}\left(1-(-1)^{p_{1}(\bar{x})}-(-1)^{p_{2}(\bar{x})}-(-1)^{p_{1}(\bar{x})+p_{2}(\bar{x})}\right)\right|.$
We will simplify these formulae using the following Lemma.
Lemma 3.
For $a,b\in\mathbf{F}_{2}$, we have
$1+(-1)^{a}+(-1)^{b}-(-1)^{a+b}=2(-1)^{ab}\,.$ (2.7)
Proof.
The proof follows easily by direct computation. We provide here, however, a complicated proof. Note that $(-1)^{x}=1-2x,x\in\mathbf{F}_{2}$. So
$(-1)^{a+b}=(-1)^{a}(-1)^{b}=(1-2a)(1-2b)=1-2(a+b)+4ab.$
So
$1+(-1)^{a}+(-1)^{b}-2(-1)^{ab}$ $=1+1-2a+1-2b-2(1-2ab)$
$=1-2(a+b)+4ab$
$=(-1)^{a+b}.\qed$
So for polynomials $p_{1},p_{2}$ taking values in $\mathbf{F}_{2}$, we have
$1+(-1)^{p_{1}}+(-1)^{p_{2}}-(-1)^{p_{1}+p_{2}}=2(-1)^{p_{1}p_{2}}\,.$
Replacing $p_{1}$ by $p_{1}+1$ and $p_{2}$ by $p_{2}+1$ yields also
$1-(-1)^{p_{1}}+(-1)^{p_{2}}+(-1)^{p_{1}+p_{2}}=2(-1)^{p_{1}p_{2}+p_{2}}\,$,
$1+(-1)^{p_{1}}-(-1)^{p_{2}}+(-1)^{p_{1}+p_{2}}=2(-1)^{p_{1}p_{2}+p_{1}}\,$,
and
$1-(-1)^{p_{1}}-(-1)^{p_{2}}-(-1)^{p_{1}+p_{2}}=2(-1)^{p_{1}p_{2}+p_{1}+p_{2}}\,.$
Using these four identities, we may rewrite the above four terms as
$I_{0}=\frac{1}{2}\sum_{\bar{A}\subset[n]\setminus[2],|\bar{A}|=k}\left|\frac{1}{2^{n-2}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x})}(-1)^{\bar{x}_{\bar{A}}}\right|,$
$I_{1}=\frac{1}{2^{n-1}}\sum_{\bar{A}\subset[n]\setminus[2],|\bar{A}|=k-1}\left|\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})+p_{2}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x})}(-1)^{\bar{x}_{\bar{A}}}\right|,$
$I_{2}=\frac{1}{2^{n-1}}\sum_{\bar{A}\subset[n]\setminus[2],|\bar{A}|=k-1}\left|\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})+p_{1}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x})}(-1)^{\bar{x}_{\bar{A}}}\right|,$
and
$I_{12}=\frac{1}{2^{n-1}}\sum_{\bar{A}\subset[n]\setminus[2],|\bar{A}|=k-2}\left|\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})+p_{1}(\bar{x})+p_{2}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x})}(-1)^{\bar{x}_{\bar{A}}}\right|.$
Note that the polynomial $p_{0}+p_{1}p_{2}$ on $\mathbf{F}_{2}^{n-2}$ is of degree at most $2$, so by definition,
$I_{0}\leq\frac{1}{2}W(2,k,n-2).$
Similarly, we have
$I_{1},I_{2}\leq\frac{1}{2}W(2,k-1,n-2),\quad I_{12}\leq\frac{1}{2}W(2,k-2,n-2).$
All combined, we have shown that
$W(2,k,n)\leq\frac{1}{2}W(2,k,n-2)+W(2,k-1,n-2)+\frac{1}{2}W(2,k-2,n-2).$
Therefore,
$W(2,k,n)$ $\leq\frac{1}{2}W(2,k,n-2)+W(2,k-1)+\frac{1}{2}W(2,k-2)$
$\leq\frac{1}{2}(1+\sqrt{2})^{k}+(1+\sqrt{2})^{k-1}+\frac{1}{2}(1+\sqrt{2})^{k-2}$
$=(1+\sqrt{2})^{k}\,.$
Here we used the induction hypotheses (2.3) and (2.4). This completes the proof.
3. Another approach via the structure theorem of Dickson
Pooya Hatami kindly explained to us what proof of the estimate of Fourier weight-level-$k$ for quadratic polynomials they had in mind in [CHHL18]. As [CHHL18] does not have the proof of this kind of estimate we include it here as it seems interesting to compare it with the proof in Section 2.
It is based on a beautiful structure theorem of Dickson (see Theorem 4 Chapter 15 of [McWS77]), which we formulate now in a convenient form. First notice that any quadratic polynomial $q:\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}$ can be written in the form
$q(x)=x^{t}Ax+(\ell,x)+c,$
where $A$ is an upper triangular matrix with $\mathbf{F}_{2}$ entries, $\ell$ is a binary vector and $c\in\mathbf{F}_{2}$.
Theorem 4 ([McWS77] and Lemma 2.4 [BGGM19]).
Let $q$ be as above and let $A$ be an upper triangular binary matrix of size $n\times n$. Let $m$ be the rank of $A+A^{t}$. Then there exists a non-singular linear mapping $T:\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}^{n}$ such that
$q\circ T^{-1}(y)=\sum_{j=1}^{m}y_{2j-1}y_{2j}+(L,y)+b,$ (3.1)
where $L\in\mathbf{F}_{2}^{n},b\in\mathbf{F}_{2}$. Moreover, if $(L,y)$ is linearly dependent on $\{y_{1},\dots,y_{2m}\}$, there exists another affine change of variable $\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}^{n}$ that transfers $q\circ T^{-1}(y)$ into $Q(z),z=(z_{1},\dots,z_{n})$, where
$Q(z)=\sum_{j=1}^{m}z_{2j-1}z_{2j}+d,$ (3.2)
where $d\in\mathbf{F}_{2}$.
Those changes of variables do not change the expectation over $\mathbf{F}_{2}^{n}$, which we will denote by $\mathbf{E}$. One simple consequence then is the fact about the Fourier coefficients of $f:=(-1)^{q(x)}$: Namely, let $S\subset[n]$, then
$|\hat{f}(S)|$ $=|\mathbf{E}(-1)^{q(x)+\sum_{k\in S}x_{k}}|$
$=\begin{cases}2^{-m},&{\sum_{k\in S}x_{k}}\in-(L,T(x))+\text{span}\{(T(x))_{j}\}_{j=1}^{2m},\\ 0,\quad&\text{otherwise}\,.\end{cases}$ (3.3)
Let us estimate the number of $S$, $|S|=k$, such that the first line holds.
3.1. Vectors of boolean weight $k$ in affine subspace of $\mathbf{F}_{2}^{n}$.
We will prove that any affine subspace of $\mathbf{F}_{2}^{n}$ of dimension $2m$ contains at most $2m+1\choose k$} vectors of boolean weight $k$.
Lemma 5.
Let $W\subset\mathbf{F}_{2}^{n}$ be an affine subspace of dimension $h$. Then $W$ contains at most $h+1\choose k$} vectors of weight $k$.
Proof.
We write $W=w+V$ for a subspace $V$ of dimension $h$. By permuting the coordinates and applying Gaussian elimination, we can choose a basis $v^{1},\dotsc,v^{h}$ of $V$ that looks as follows:
$\begin{pmatrix}--v^{1}--\\ --v^{2}--\\ \vdots\\ --v^{h}--\end{pmatrix}=\begin{pmatrix}1&0&\ldots&0&*&*&\dotsb&*\\ 0&1&\ldots&0&*&*&\dotsb&*\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&&\vdots\\ 0&0&\ldots&1&*&*&\dotsb&*\end{pmatrix}\,$.
Now we can further assume that $w_{1}=\dotsb=w_{h}=0$. We will prove the lemma by induction.
The lemma holds when $k=1$. Indeed, any sum of at least $2$ vectors $v^{j}$ is nonzero in at least two of the coordinates $1,\dotsc,h$, and hence its sum with $w$ does not have weight $1$. So there are at most $h+1$ vectors in $w+V$ that could have weight $1$, namely
$w,w+v^{1},\dotsc,w+v^{h}\,.$
Now let us assume that it holds for k1k-1. If the coefficient of $v^{1}$ in a weight $k$-vector $v\in w+V$ is nonzero, then we can write
$v=\delta^{1}+(w+v^{1}-\delta^{1})+r$
where $r\in V^{\prime}:=\text{span}\{v^{2},\dotsc v^{h}\}$ and $\delta^{1}$ is the vector that is one in coordinate 11 and zero everyhwere else. Then $(w+v^{1}-\delta^{1})+r$ is a weight (k1)(k-1) vector in an (h1)(h-1)-dimensional affine space. Moreover, $w^{\prime}=(w+v^{1}-\delta^{1})$ satisfies $w^{\prime}_{i}=0$ for $i=1,\dotsc,h$. By induction hypothesis, there are thus at most
$h\choose k-1$}
choices for $v$ with nonzero coefficient of $v^{1}$.
On the other hand, if the coefficient of $v^{1}$ vanishes, then $v$ itself is a weight $k$ vector in the dimension h1h-1 affine subspace $w+V^{\prime}=w+\langle v^{2},\dotsc,v^{h}\rangle$. By induction again, there are at most
$h\choose k$}
such vectors $v$. Thus, in total there are at most
${h\choose k-1}+{h\choose k}={h+1\choose k}$
vectors $v$. This completes the proof. ∎
The estimate in Lemma 5 is sharp: If $W$ is the affine subspace in $\mathbf{F}_{2}^{n}$ generated by
$\{x\in\mathbf{F}_{2}^{n}\,:\,\sum_{i=1}^{n}x_{i}=k\}$
then the dimension of $W$ is n1n-1, and $W$ contains exactly
${n\choose k}={\dim W+1\choose k}$
vectors of weight $k$, namely all weight $k$ vectors in $\mathbf{F}_{2}^{n}$.
3.2. Estimating the Fourier weight
Combining (3.3) and Lemma 5 shows that
$\sum_{|S|=k}|\hat{f}(S)|\leq 2^{-m}{2m+1\choose k}\,$.
To further estimate this we start with:
Lemma 6.
For all $\alpha\in[0,1]$, we have
$\alpha^{\alpha}(2-\alpha)^{2-\alpha}\geq 2(\sqrt{2}-1)^{\alpha}\,.$
Proof.
Substituting $\alpha=2(1-y)$ turns this into the equivalent inequality
$\left(\frac{y}{2^{-1/2}}\right)^{y}\left(\frac{1-y}{1-2^{-1/2}}\right)^{1-y}\geq 1\,.$
It is readily verified that equality holds at $y=2^{-1/2}$. Furthermore, if $h(y)=\exp(g(y))$ denotes the left hand side of the inequality, then
$\frac{\mathrm{d}}{\mathrm{d}y}h(y)=\exp(g(y))g^{\prime}(y)=\exp(g(y))\left(\log\frac{y}{2^{-1/2}}-\log\frac{1-y}{1-2^{-1/2}}\right)\,.$
Thus, the derivative vanishes at $y=2^{-1/2}$. Finally, we have
$\frac{\mathrm{d}^{2}}{\mathrm{d}y^{2}}h(y)=\exp(g(y))((g^{\prime}(y))^{2}+g^{\prime\prime}(y))\geq\exp(g(y))g^{\prime\prime}(y)$
$\exp(g(y))\left(\frac{1}{y}+\frac{1}{1-y}\right)\,$,
so $h$ is convex and the inequality follows. ∎
Lemma 7.
For all $m\geq 1$ and $0\leq k\leq m$
${2m+1\choose k}\leq\sqrt{\frac{2}{\pi}}2^{m}k^{-1/2}(1+\sqrt{2})^{k}\,.$
This is sharp in the sense that there exist a sequence of $(m,k)$ such that the ratio of left hand side and right hand side tends to $1$.
Proof.
We use the estimate (see e.g. [McWS77], page 309)
${n\choose k}\leq\sqrt{\frac{n}{2\pi k(n-k)}}2^{nH(k/n)}$ (3.4)
where $H$ is the binary entropy function
$H(p)=-p\log_{2}p-(1-p)\log_{2}(1-p).$
Note that by Stirling’s formula for $k/n$ fixed and $n\to\infty$ the bound (3.4) is sharp, in the sense that the ratio of the left hand side and the right hand side approaches one.
Using (3.4), we obtain for $k=\alpha(m+1/2)$:
$2^{-m}{2m+1\choose k}$ $\leq\frac{1}{\sqrt{\pi k(2-\alpha)}}2^{(2m+1)H(\alpha/2)-m}$
$=\sqrt{\frac{2}{\pi k(2-\alpha)}}\left(\frac{2}{\alpha^{\alpha}(2-\alpha)^{(2-\alpha)}}\right)^{m+1/2}\,.$ (3.5)
Bounding the second factor with Lemma 6 and the first factor with $\alpha<1$
$\leq\sqrt{\frac{2}{\pi}}k^{-1/2}(1+\sqrt{2})^{k}\,.$
For sharpness we use the following claim: There exist infinitely many pairs of integers $(k,m)$ such that
$\big|\frac{k}{2m+1}-\frac{2-\sqrt{2}}{2}\big|\leq\frac{1}{m^{2}}\qquad\iff\qquad|\alpha-(2-\sqrt{2})|\leq\frac{2}{m^{2}}.$ (3.6)
Using this, the fact that (3.4) and therefore (3.5) are asymptotically sharp, and that equality in Lemma 6 holds at $2-\sqrt{2}$ yields the claimed sharpness.
The claim (3.6) holds since for any convergent $p/q$ in the continued fraction expansion of $1-\sqrt{2}/2$ we have
$\left|\frac{p}{q}-\frac{2-\sqrt{2}}{2}\right|\leq\frac{1}{q^{2}}.$
Further, two consecutive convergents $p/q$ and $p^{\prime}/q^{\prime}$ satisfy
$pq^{\prime}-qp^{\prime}=\pm 1,$
so that at least one of $q,q^{\prime}$ is odd. See [HW08, Theorem 171] and related discussion for the proof. ∎
We got the desired estimate for $k\leq m$. If $m+1\leq k\leq 2m+1$ we just put $\ell=2m+1-k<k$ and get the previous estimate with $\ell$ instead of $k$. We are done:
$W(2,k,n)\leq\sqrt{\frac{2}{\pi}}k^{-1/2}(1+\sqrt{2})^{k},$
where we used that function $x\to\frac{(1+\sqrt{2})^{x}}{x^{1/2}}$ increases on $1,\infty)$.
3.3. Sharpness
In fact, the arguments in this section were all sharp: Let
$q(x)=\sum_{j=1}^{m}x_{2j}x_{2j-1}$
and pick $(m,k)$ an asymptotically sharp sequence for Lemma 7, so
$k\approx(2-\sqrt{2})m\,.$
We use (3.3) to obtain
$\sum_{|S|=k}|\hat{f}(S)|=2^{-m}{2m\choose k}=2^{-m}{2m+1\choose k}\frac{2m+1-k}{2m+1}.$
By the choice of $k$ we are in the essential equality case of Lemma 7, which yields
$\sum_{|S|=k}|\hat{f}(S)|\sim\sqrt{\frac{2}{\pi}}k^{-1/2}(1+\sqrt{2})^{k}\,.$
Thus the bound in Theorem 2 is sharp.
4. Degree 3 case
Here are some thoughts on the degree 3 case. Suppose that $\deg(p)=3$ and without loss of generality
$p(x)=$ $x_{1}x_{2}x_{3}+x_{1}p_{1}(\bar{x})+x_{2}p_{2}(\bar{x})+x_{3}p_{3}(\bar{x})$
$\ \ \ +x_{1}x_{2}p_{12}(\bar{x})+x_{1}x_{3}p_{13}(\bar{x})+x_{2}x_{3}p_{23}(\bar{x})+p_{0}(\bar{x})$
where $x=(x_{1},x_{2},x_{3},\bar{x})\in\mathbf{F}_{2}^{n}$. Then
$\sum_{A\subset[n],|A|=k}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)+x_{A}}\right|=I_{0}+I_{1}+I_{2}+I_{3}+I_{12}+I_{13}+I_{23}+I_{123},$
where
$I_{0}=\sum_{A\cap[3]=\emptyset,|A|=k}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)+x_{A}}\right|$ (4.1)
and for $\emptyset\neq J\subset[3]$
$I_{J}=\sum_{A\cap[3]=\{J\},|A|=k}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)+x_{A}}\right|.$ (4.2)
Denoting $4,n]:=[n]\setminus[3]$ and $\bar{A}:=A\setminus[3]$. By definition,
$I_{0}$ $=\sum_{A\cap[3]=\emptyset,|A|=k}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)+x_{A}}\right|$
$=\sum_{\bar{A}\subset[4,n],|\bar{A}|=k}\left|\frac{1}{2^{n}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-3}}(-1)^{\bar{x}_{\bar{A}}}\sum_{x_{1},x_{2},x_{3}\in\mathbf{F}_{2}}(-1)^{p(x)}\right|.$
By definition,
$p(0,0,0,\bar{x})=p_{0}(\bar{x}),$
$p(1,0,0,\bar{x})=p_{1}(\bar{x})+p_{0}(\bar{x})$
$p(0,1,0,\bar{x})=p_{2}(\bar{x})+p_{0}(\bar{x})$
$p(0,0,1,\bar{x})=p_{3}(\bar{x})+p_{0}(\bar{x})$
$p(1,1,0,\bar{x})=p_{1}(\bar{x})+p_{2}(\bar{x})+p_{12}(\bar{x})+p_{0}(\bar{x})$
$p(1,0,1,\bar{x})=p_{1}(\bar{x})+p_{3}(\bar{x})+p_{13}(\bar{x})+p_{0}(\bar{x})$
$p(0,1,1,\bar{x})=p_{2}(\bar{x})+p_{3}(\bar{x})+p_{23}(\bar{x})+p_{0}(\bar{x})$
$p(1,1,1,\bar{x})=1+p_{1}(\bar{x})+p_{2}(\bar{x})+p_{3}(\bar{x})+p_{12}(\bar{x})+p_{13}(\bar{x})+p_{23}(\bar{x})+p_{0}(\bar{x}).$
Thus,
$\sum_{x_{1},x_{2},x_{3}\in\mathbf{F}_{2}}(-1)^{p(x)}$ $=(-1)^{p_{0}(\bar{x})}\left[1+(-1)^{p_{1}(\bar{x})}+(-1)^{p_{2}(\bar{x})}+(-1)^{p_{3}(\bar{x})}\right$.
$+\left.(-1)^{p_{1}(\bar{x})+p_{2}(\bar{x})+p_{12}(\bar{x})}+(-1)^{p_{1}(\bar{x})+p_{3}(\bar{x})+p_{13}(\bar{x})}+(-1)^{p_{2}(\bar{x})+p_{3}(\bar{x})+p_{23}(\bar{x})}\right$.
$+\left.(-1)^{1+p_{1}(\bar{x})+p_{2}(\bar{x})+p_{3}(\bar{x})+p_{12}(\bar{x})+p_{13}(\bar{x})+p_{23}(\bar{x})}\right].$ (4.3)
We have the following lemma.
Lemma 8.
We have the following identity: for any $a,b,c,x,y,z\in\mathbf{F}_{2}$
$1+(-1)^{a}+(-1)^{b}+(-1)^{c}+(-1)^{a+b+z}+(-1)^{a+y+c}+(-1)^{x+b+c}$
$+(-1)^{1+a+b+c+x+y+z}$
$=$ $2(-1)^{ab}+2(-1)^{ab+ax+by+xy+c}+2(-1)^{a+b+xz+yz+cz}.$
Proof.
Recall that we have for all $\alpha,\beta\in\mathbf{F}_{2}$
$1+(-1)^{\alpha}+(-1)^{\beta}+(-1)^{1+\alpha+\beta}=2(-1)^{\alpha\beta}.$ (4.4)
Then we have
$1+(-1)^{a}+(-1)^{b}=2(-1)^{ab}-(-1)^{1+a+b},$ (4.5)
$(-1)^{c}+(-1)^{a+y+c}+(-1)^{x+b+c}$ $=(-1)^{c}\left[1+(-1)^{a+y}+(-1)^{x+b}\right]$
$=(-1)^{c}\left[2(-1)^{(a+y)(x+b)}-(-1)^{1+a+b+x+y}\right]$
$=2(-1)^{c+ab+ax+by+xy}-(-1)^{1+a+b+c+x+y},$
and
$(-1)^{a+b+z}+(-1)^{1+a+b+c+x+y+z}$ $=(-1)^{a+b}[(-1)^{z}+(-1)^{1+c+x+y+z}]$
$=(-1)^{a+b}\left[(-1)^{z}+(-1)^{1+c+x+y+z}\right]$
$=(-1)^{a+b}\left[2(-1)^{(c+x+y)z}-(-1)^{c+x+y}-1\right]$
$=2(-1)^{a+b+cz+xz+yz}-(-1)^{a+b+c+x+y}-(-1)^{a+b}.$
Summing all these identities finishes the proof. ∎
According to this lemma, we may rewrite
$\sum_{x_{1},x_{2},x_{3}\in\mathbf{F}_{2}}(-1)^{p}=$ $2(-1)^{p_{0}+p_{1}p_{2}}+2(-1)^{p_{0}+p_{1}p_{2}+p_{1}p_{23}+p_{2}p_{13}+p_{13}p_{23}+p_{3}}$
$+2(-1)^{p_{0}+p_{1}+p_{2}+p_{12}p_{23}+p_{12}p_{13}+p_{3}p_{12}}.$ (4.6)
This allows us to reduce eight terms into six terms. Recall that $\deg(p_{i})\leq 2$ and $\deg(p_{ij})\leq 1$, so we see that most new polynomials have degrees at most $3$, except for $p_{1}p_{2}$. Can we solve this problem by rewriting the identity in the above lemma?
The above discussion for $I_{0}$ applies to $I_{J},J\neq\emptyset$ in a similar way. Take $I_{1}$ for example. By definition,
$I_{1}$ $=\sum_{A\cap[3]=\{1\},|A|=k}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)+x_{A}}\right|$
$=\sum_{\bar{A}\subset[4,n],|\bar{A}|=k-1}\left|\frac{1}{2^{n}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-3}}(-1)^{\bar{x}_{\bar{A}}}\sum_{x_{1},x_{2},x_{3}\in\mathbf{F}_{2}}(-1)^{p(x)+x_{1}}\right|.$
So, to rewrite the sum of powers of 1-1, we only need to replace $p$ in $I_{0}$ with $p+x_{1}$. Recalling the definition of $p$, we only need to replace $p_{1}$ with $p_{1}+1$. That is,
$\sum_{x_{1},x_{2},x_{3}\in\mathbf{F}_{2}}(-1)^{p+x_{1}}=$ $2(-1)^{(p_{1}+1)p_{2}}+2(-1)^{(p_{1}+1)p_{2}+(p_{1}+1)p_{23}+p_{2}p_{13}+p_{13}p_{23}+p_{3}}$
$+2(-1)^{(p_{1}+1)+p_{2}+p_{12}p_{23}+p_{12}p_{13}+p_{3}p_{12}}.$
Again, the main issue is the degree of $p_{1}p_{2}$.
In general, for $I_{J}$ with $\emptyset\neq J\subset[3]$, the difference is that one replaces all $p_{i},i\in J$ with $p_{i}+1$.
5. Other arguments
The other argument to deal with degree $3$ by using the identity
$(-1)^{\alpha}+(-1)^{\beta}+(-1)^{\gamma}-(-1)^{\alpha+\beta+\gamma}=2(-1)^{\alpha\beta+\alpha\gamma+\beta\gamma}$
leads also to higher degree polynomial terms, and we have no similar identity for degree $4$, so we cannot iterate there.
However, one can prove the following theorems.
Theorem 9.
Suppose that polynomial $p$ is such that $\deg p=3$ but $\deg p_{0}=2$. Then the $k$-th Fourier weight of $(-1)^{p}$ is at most $C(1+\sqrt{2})^{k}$ for all $k$.
This theorem can be used by tracking the consequences of (4) if $\deg p_{0}\leq 2$. In fact, the assumption says that
$p=x_{1}\cdot p_{1}+x_{2}\cdot p_{2}+x_{3}\cdot p_{3}+p_{0},$
where all $p_{i},i=0,1,2,3$ are degree $2$ polynomials. The same simple tracking of (4) will work for any constant $L$ independent of $n$ if $p$ is of degree $2$ in $x_{L+1},...,x_{n}$. So this theorem is a very modest improvement.
The next result is also based on the decomposition of degree $3$ polynomial $p$ into a restricted number of products of linear and quadratic polynomials plus (very roughly speaking) another polynomial of degree $3$ of much smaller number of variables. The next result uses the definition of bias of $p$, see e.g. [HS09].
Bias of $p:\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}$ is
$\text{bias}(p)=\big|\mathbf{E}(-1)^{p}\big|\,.$
Not too close to zero bias means that the distribution created by $p$ is not too close to the uniform distribution. Random polynomials have small bias. So the assumption is that $p$ is far from being random. Kaufman and Lovett [KL08] (see also [GT08] for a close but different situation) proved that polynomial on $\mathbf{F}_{2}^{n}$ with very large $n$ and of degree $d$ with noticeable bias can be represented as a certain function of smaller degrees whose number depends on $c=c(d,\text{bias})$.
Theorem 10.
Suppose that polynomial $p$ is such that $\deg p=3$ but $\text{bias}(p)\geq\delta$. Then the $k$-th Fourier weight of $(-1)^{p}$ is at most $(C/\delta)^{k}$ for all $k$.
The proof of this theorem in [HS09] is based on a structure theorem (not unlike Dickson’s lemma used above) that claims that a polynomial of degree $3$ with a noticeable bias $\delta$ has a decomposition to the sum of products of linear and quadratic polynomials, where the sum has at most $c\log\frac{1}{\delta}$ terms plus a cubic polynomial of certain amount of linear terms, whose number is again controlled by $\delta$ (and independent of $n$).
6. Why the induction argument for $k>2$ does not work
We use three things:
$W(d,1,n)\leq 4d\leq 2^{3d}.$ (6.1)
$W(2,k,n)\leq C_{0}^{k}2^{3\cdot 2}\leq(C_{0}2^{3\cdot d}+1)^{k},\quad d=2\,.$ (6.2)
and Lemma 6.2 of [CHHL18]:
Lemma 11 (CHHL).
$W(d,k,n)^{2}\leq 2^{2k}W(d-1,2k,n)+W(d,k,n)\cdot\sum_{\ell=1}^{k}{k\choose\ell}W(d,k-\ell,n)\,.$
We would like to use the above lemma, along with(6.1) and (6.2) to prove by induction on $d$ and $k$ the estimate
$W(d,k,n)\leq(C_{0}\cdot 2^{3\cdot d}+1)^{k}\,.$ (6.3)
By (6.1) and (6.2) the estimate holds for $d=1,2$ and all $k$. Our goal is to show it for the pair $(d,k)$, and we assume by induction that it holds for all pairs $(d^{\prime},k^{\prime})$ such that $d<d^{\prime}$ or such that $d=d^{\prime}$ and $k^{\prime}<k$. By Lemma 11 and the induction hypothesis, we have
$W(d,k,n)^{2}$ $\leq 2^{2k}W(d-1,2k,n)+W(d,k,n)\cdot\sum_{l=1}^{k}{k\choose l}W(d,k-l,n)$
$\leq 2^{2k}(C_{0}\cdot 2^{3\cdot(d-1)}+1)^{2k}+W(d,k,n)\cdot\sum_{l=1}^{k}{k\choose l}(C_{0}2^{3d}+1)^{k-l}$
$=2^{2k}(C_{0}\cdot 2^{3\cdot(d-1)}+1)^{2k}+W(d,k,n)((C_{0}\cdot 2^{3\cdot d}+2)^{k}-(C_{0}\cdot 2^{3\cdot d}+1)^{k})\,.$
Again, we would like to show that $W(d,k)\leq(C_{0}\cdot 2^{3\dot{d}}+1)^{k}$. Suppose that it were bigger, i.e.
$W(d,k,n)>(C_{0}\cdot 2^{3\dot{d}}+1)^{k},$
then we can divide by $W(d,k)$ and obtain
$W(d,k,n)\leq\left(\frac{2(C_{0}2^{3(d-1)}+1)}{C_{0}2^{3d}+1}\right)^{k}+((C_{0}2^{3d}+2)^{k}-(C_{0}2^{3d}+1)^{k})$ (6.4)
We have
$\frac{2(C_{0}2^{3(d-1)}+1)}{C_{0}2^{3d}+1}=\frac{\frac{1}{4}C_{0}2^{3d}+2}{C_{0}2^{3d}+1}\leq\frac{C_{0}2^{3d}+1}{C_{0}2^{3d}+1}=1\,.$
We want to show that the second term in (6.4) becomes much larger than $(C_{0}\cdot 2^{3\cdot d}+1)^{k}$ when $k$ is very large. We have for general $A$ that
$(A+1)^{k}-A^{k}=A^{k}((1+\frac{1}{A})^{k}-1)\geq A^{k}\frac{k}{A}\,.$
Using this for $A=C_{0}\cdot 2^{3\cdot d}+1$, we see that when $k>C_{0}\cdot 2^{3\cdot d}+1$, then
$(C_{0}2^{3d}+2)^{k}-(C_{0}2^{3d}+1)^{k}>(C_{0}2^{3d}+1)^{k}\,$,
so we do not get the contradiction we would like to have.
However, for $k\leq 3d$ we have ($A=C_{0}\cdot 2^{3\cdot d}+1$)
$1+(A+1)^{k}-A^{k}=1+A^{k}((1+\frac{1}{A})^{k}-1)\leq 1+A^{k}\frac{2^{k}}{A}\leq A^{k}\,.$
The estimate works as long as $k\ll 2^{d}$, but this is not helpful, because in the induction argument we pass to much larger values of $k$ than what we started with. (In fact, we have to use the lemma d2d-2 times to get to the bound for $d=2$, so the largest $k$ we get is exactly $2^{d-2}$ times the $k$ we started with.
However, this is what we get as the estimate of $W(d,k,n)$ for a certain range of $d,k$:
$W(d,k,n)\leq(C_{0}2^{3d}+1)^{k}\leq(2C_{0}2^{3d})^{k}\quad\,\,\text{if}\,\,{k\leq 3d}\,.$ (6.5)